You want to have a 70,000 college fund in 16years how much would I have invested at 5%?
Answer by stanbon(75887)

You want to have a 70,000 college fund in 16years
how much would I have invested at 5% ?
----
Assuming simple interest::
A(t) = P(1+rt)
70,000 = P(1 + 0.05*16)
70,000 = P(1.8)
P = $38,888.88
---------------
Cheers,
Stan H.
-------------
-------